OpenStudy (jazzyluvsu123):

what percent do you think out of this world people gets bullied I think 99% because I get bullied it hurts so wat percent do you think it is?

OpenStudy (anonymous):

This isnt a place for this please leave

OpenStudy (doc.brown):

\[\frac{99}{100}\times 7.12\times 10^{9}=7.05\times 10^{9}\] Just to clarify: You think a little over seven billion people get bullied because you get bullied, and it hurts. This is not a logical argument.

OpenStudy (jazzyluvsu123):

ok ?